OpenStudy (anonymous):

Check my answer? Write an equation for the given function given the period, phase shift, and vertical shift. tangent function, period =1/3 π phase shift =-1/4 π , vertical shift = -5. y=tan[3θ - 3/4π] -5

jimthompson5910 (jim_thompson5910):

I think the phase shift is positive, but everything else looks correct

OpenStudy (anonymous):

I was thinking it would be but I just wasn't sure c: thank you n.n

jimthompson5910 (jim_thompson5910):

you're welcome I'm reading that the general form is y = A*tan(Bx - C) + D where the phase shift is p = C/B, so that's why I'm thinking it's positive

OpenStudy (anonymous):

But wouldn't the negative override the positive making the phase shift negative?

jimthompson5910 (jim_thompson5910):

well y=tan[3θ - 3/4π] -5 means that B = 3 and C = 3/4π So p = C/B p = (3/4π)/3 p = π/4

OpenStudy (anonymous):

ohhh okai I see what you mean